OpenStudy (anonymous):

what is y=2x+2 in point slop form

alones (alones):

https://www.desmos.com/ use it post your equationt there and see if it works

OpenStudy (anonymous):

dident work

alones (alones):

Hrmm let me try it for you

alones (alones):

http://prntscr.com/aat5a0

OpenStudy (anonymous):

ok but i need help with wre writing the equation into point slope form

alones (alones):

Oh you mean this y=2x+2 Step 1: Flip the equation. 2x+2=y Step 2: Add -2 to both sides. 2x+2+−2=y+−2 2x=y−2 Step 3: Divide both sides by 2. 2x/2=y−2/2 x=1/2y−1

OpenStudy (anonymous):

yes